OpenStudy (anonymous):

Part 1: [4 points] Find the measures of central tendency (mean, median, mode and range) for the following data set: 14, 27, 24, 29, 28, 25, 14, 26 Part 2: [3 points] Tell which of these measures would best represent the whole data set and why.

OpenStudy (anonymous):

get them in numerical order first

OpenStudy (anonymous):

14,14,24,25,26, 27,28,29

OpenStudy (anonymous):

Ok .

OpenStudy (anonymous):

Mean= average. add together all the values, and divide by the number of values. 14+14+24...=187

OpenStudy (anonymous):

since there are 8 numbers, divide 187 by 8 23.375 mean=23.375

OpenStudy (anonymous):

what about the median and range and mode ?

OpenStudy (anonymous):

median=middle number to find the number in the middle, start with one finger at each end and begin crossing numbers off. for every number you cross of on the right side, you should do one on the left, and that will eventually bring you to the middle number.

OpenStudy (anonymous):

ur left with 25,26 in the middle. find the average of those two. 25+26/2=25.5 median=25.5

OpenStudy (anonymous):

mode=most. which number occurs the most times in this set? answer=14 mode=14

OpenStudy (anonymous):

range ?

OpenStudy (anonymous):

Range= biggest number in ur set minus the smallest number 29-14 15 Range=15

OpenStudy (anonymous):

Okay what about for part 2 ?

OpenStudy (anonymous):

mean=23.375 median=25.5 mode=14 range=15

OpenStudy (anonymous):

which number best represents the whole set?

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

Not the mode or the range, because they're way too low

OpenStudy (anonymous):

so the median ?

OpenStudy (anonymous):

so you're stuck between the median and the mean. Use your own judgement

OpenStudy (anonymous):

do i add ? or multiply ?